OpenStudy (anonymous):

find the value of the limit for the sequence given {1.9.17...(8n+1)/(8n)^2}

OpenStudy (anonymous):

@jim_thompson5910 sorry to bother you again but for this problem we just focus on (8n+1)/(8n)^2 right?

jimthompson5910 (jim_thompson5910):

why are there multiple decimal points in 1.9.17

jimthompson5910 (jim_thompson5910):

that's just confusing notation

jimthompson5910 (jim_thompson5910):

are they supposed to be commas?

OpenStudy (anonymous):

idk no they are points that is how my teacher gave us our homework

jimthompson5910 (jim_thompson5910):

are you able to take a screenshot and show me the full problem?

OpenStudy (anonymous):

yeah i'll do that

OpenStudy (anonymous):

here it is

jimthompson5910 (jim_thompson5910):

ok let me look it over

jimthompson5910 (jim_thompson5910):

ok let's say n = 1. What is the value of that expression you gave in the screenshot?

jimthompson5910 (jim_thompson5910):

you can give an approximation

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

my bad its .1406

jimthompson5910 (jim_thompson5910):

8n+1 = 9 when n = 1 so the numerator will have 1*9 = 9 the denominator is equal to (8n)^2 = (8*1)^2 = 64 So the expression is equal to 9/64 = 0.140625

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

when n = 2, what is the expression equal to?

OpenStudy (anonymous):

.0664

OpenStudy (anonymous):

its approaching zero isnt it?

jimthompson5910 (jim_thompson5910):

8n+1 = 17 when n = 2 so the numerator will have 1*9*17 = 153 the denominator is equal to (8n)^2 = (8*2)^2 = 256 So the expression is equal to 153/256 = 0.59765625

jimthompson5910 (jim_thompson5910):

the numerator is NOT simply 17 it's 17 times the previous values (1 and 9)

jimthompson5910 (jim_thompson5910):

when n = 3, the numerator isn't simply 25 it is 25*17*9*1 = 3825

OpenStudy (anonymous):

oh so we have to multiply times the other three values

jimthompson5910 (jim_thompson5910):

the numerator is basically the product function \[\Large \prod_{k=0}^{n} (8n+1) \]

jimthompson5910 (jim_thompson5910):

The big upper case pi stands for product

jimthompson5910 (jim_thompson5910):

It's analogous to the sigma sum function

jimthompson5910 (jim_thompson5910):

oops typo, meant to write \[\Large \prod_{k=0}^{n} (8k+1)\]

jimthompson5910 (jim_thompson5910):

\[\Large \prod_{k=0}^{1} (8k+1) = 1*9\] \[\Large \prod_{k=0}^{2} (8k+1) = 1*9*17\] \[\Large \prod_{k=0}^{3} (8k+1) = 1*9*17*25\] ... ... ... \[\Large \prod_{k=0}^{n} (8k+1) = 1*9*17*25*\ldots*(8n+1)\]

OpenStudy (anonymous):

oh so we are just taking the limit of the denominator?

OpenStudy (anonymous):

or the whole thing

jimthompson5910 (jim_thompson5910):

we're doing the whole thing. It's a bit tedious, so I had excel do much of the work check out the attached document

jimthompson5910 (jim_thompson5910):

notice at n = 26, we have the whole expression equal to 4.51 * 10^45 which is a very huge number

OpenStudy (anonymous):

oh man did you create that chart?

OpenStudy (anonymous):

those are some pretty huge numbers man

jimthompson5910 (jim_thompson5910):

I used the computer to do much of the work

OpenStudy (anonymous):

thats pretty cool man

OpenStudy (anonymous):

so the value for this limit problem is infinity

jimthompson5910 (jim_thompson5910):

yes